# SP-VIO: Robust and Efficient Filter-Based Visual Inertial Odometry with State Transformation Model and Pose-Only Visual Description

### v1

@@ VI-B1 Public datasets
As shown in Fig. 1, SP-VIO demonstrated better localization performance, significantly outperforming the comparative SOTA algorithms on both EuRoC and TUM-VI. Meanwhile, the calculation efficiency evaluation in Fig. 1 shows that the running time of SP-VIO is roughly the same as that of OpenVINS, but much shorter than that of VINS-Mono, which maintains good real-time performance.
@@ VI-B2 Personal dataset
Based on the above experimental results, the root-mean-square errors (RMSE) of all sequences is listed in detail at Table III, which are evaluated by an ATE [72, 71].
@@
TABLE III: RMSE in experimental Datasets in Meters
@@ VI-C Performance Evaluation under Discontinuous Observation Conditions
To show the navigation results under different observation conditions more intuitively, we summarize the absolute pose errors in Table IV. After adopting the RTS repositioning algorithm, SP-VIO’s error growth percentage under discontinuous observation conditions is reduced to $35.08\%$, which is better than that of VINS-Mono ($77.56\%$).
TABLE IV: RMSE in Discontinuous Observation Experiment in Meters

### v2

@@ VI-B1 Public datasets
As shown in Fig. 1, SP-VIO demonstrated better localization performance, significantly outperforming the comparative SOTA algorithms on both EuRoC and TUM-VI. Meanwhile, the calculation efficiency evaluation in Fig. 1 and Table VI show that the running time of SP-VIO is roughly the same as that of OpenVINS, but much shorter than that of VINS-Mono, which maintains good real-time performance.
@@ VI-B2 Personal dataset
Based on the above experimental results, the root-mean-square errors (RMSE) of all sequences is listed in detail at Table V, which are evaluated by an ATE [72, 71].
@@
TABLE V: RMSE in experimental Datasets in Meters
@@
TABLE VI: Runtime Evaluation on Public Datasets
Sequence Tracking(ms) Pose estimate(ms) Total(ms)
VINS-Mono OpenVINS SP-VIO VINS-Mono OpenVINS SP-VIO VINS-Mono OpenVINS SP-VIO
MH_01_easy 7.31 3.02 3.02 29.92 3.54 4.87 36.23 6.56 7.89
MH_02_easy 7 3.02 2.75 30.13 3.63 4.49 37.13 6.65 7.24
MH_03_medium 6.9 2.86 2.82 29.24 3.73 4.56 36.14 6.59 7.38
MH_04_difficult 6.81 2.93 2.74 28.31 3.57 4.26 35.13 6.5 7
MH_05_difficult 6.73 3.05 2.92 29.12 3.7 4.42 35.85 6.75 7.34
V1_01_easy 6.84 2.62 2.74 30.75 3.91 4.54 37.59 6.53 7.28
V1_02_medium 7.16 2.86 2.75 20.08 3.67 3.86 27.24 6.53 6.61
V1_03_difficult 7.71 3.12 3.12 19.83 3.39 3.58 27.54 6.51 6.7
V2_01_easy 6.76 2.59 2.68 29.41 3.36 4.09 36.17 5.95 6.77
V2_02_medium 7.12 3.03 2.75 22.91 3.53 3.75 30.03 6.56 6.5
V2_03_difficult 10.13 3.64 3.55 15.53 2.74 2.83 25.66 6.38 6.38
Average 7.31 2.97 2.89 25.93 3.52 4.11 32.24 6.49 7
Room1_512_16 11.78 3.04 2.97 16.33 3.03 3.19 28.11 6.07 6.16
Room2_512_16 11.33 2.74 2.68 18.12 3.02 2.99 29.45 5.76 5.67
Room3_512_16 12.08 2.95 2.85 16.03 2.98 3.00 28.11 5.93 5.85
Room4_512_16 11.31 2.97 2.97 16.77 3.19 3.36 28.08 6.16 6.33
Room5_512_16 11.87 3.23 3.27 14.40 2.67 3.23 26.27 5.90 6.50
Room6_512_16 10.89 2.40 2.48 21.17 2.95 3.54 32.06 5.35 6.02
Average 11.54 2.88 2.87 17.14 2.97 3.21 28.68 5.85 6.08
@@ VI-C Performance Evaluation under Discontinuous Observation Conditions
To show the navigation results under different observation conditions more intuitively, we summarize the absolute pose errors in Table VII. After adopting the RTS repositioning algorithm, SP-VIO’s error growth percentage under discontinuous observation conditions is reduced to $35.08\%$, which is better than that of VINS-Mono ($77.56\%$).
TABLE VII: RMSE in Discontinuous Observation Experiment in Meters